Question 8

10 / 10 pts

Congratulations! You just purchased a new Malibu Wake surf boat. The sales price of the boat was $118,000. You financed 95% of the sales price at 5.375% over 12 years. What is your monthly payment?

Question 9

5 / 5 pts

Based on the information from the previous question, how much of your first payment goes to interest?

IncorrectQuestion 10

0 / 5 pts

Based on the boat purchase, if you make payments for 12 years, how much total interest will you pay?

IncorrectQuestion 11

0 / 10 pts

You are thinking about purchasing a new John Deere S790 combine. The purchase price is $585,000. You plan to operate the combine for 3 years. You predict that you can sell it for $395,000 at the end of 3 years. What is your annualized cost of owning this combine if your required return on capital is 7.25%?
Answer #1

8)

time period =12 x 12 = 144 months

monthly rate (i) = 5.375%/12 = 0.448%

Loan amount = 95% x 118000 = 112100

monthly payment = m

112100 = m/(1 + 0.448%) + m/(1 + 0.448%)2 +..... + m/(1 + 0.448%)144

m = $ 1058

9)

Interest payment (first payment) = 0.448% x 112100 = 502.11

10)

Total interest = 1058 x 144 - 112100 = 40252

11)

PV = -585000 + 395000/(1 + 7.25%)3 = - 264811.9

annualized cost = A


A/(1 + 7.25%) + A/(1 + 7.25%)2 + A/(1 + 7.25%)3 = 264811.9

A = 101368.25